OpenStudy (anonymous):

Eric knew his trip to Lawrenceburg was 548 miles. If he drove for 8.9 hours, what mph did he average on his trip? You may use the formula d = rt. Round your answer to the nearest whole number

OpenStudy (anonymous):

d=vt v=d/t=548/8.9=61.573... which rounds to 62mph